# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

### v1

@@ 4.3 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$
While our main objective in this paper is to design a fast and accurate tuning strategy for Lasso, some intermediate products of our algorithm show promising empirical results in statistical tasks relevant in high-dimensional regression. Here we illustrate some preliminary results for two such tasks, estimation of noise variance $\sigma^{2}$, and a visual diagnostic like screeplot to check the sparsity assumption.
@@ 4.3.1 Noise scale estimation for high-dimensional inference
We empirically contrast $\sigma$ estimation of $\mathsf{autotune}$ against Scaled Lasso, Organic Lasso and Natural Lasso [yu2019estimating]. We also investigate the quality of estimated $\sigma$ of different algorithms by using it in inference of high-dimensional regression coefficients via $\mathsf{hdi}$ R package [dezeure2015high]. We report the sensitivity of $\sigma$ estimation to chosen $\alpha$ in the $\mathsf{autotune}$’s sequential F-test (see Sigma-Update) in the Sec.˜A.5.
@@
We at first compute the difference between algorithm’s estimated $\sigma^{2}$ and the best empirical estimate of noise level, $\sigma^{2}_{\mathrm{empirical}}=\mathrm{Var}(\text{Y}-\text{X}\beta_{\mathrm{true}})$ i.e., variance of the residuals obtained via regression of Y on X using true coefficients. Then we plot the boxplots of the difference $\widehat{\sigma}^{2}-\sigma^{2}_{\mathrm{empirical}}$ across 100 replications in Fig.˜11.
For inference, we use De-Sparsified Lasso [dezeure2015high] from $\mathsf{hdi}$ R package. It computes the Lasso coefficients for the dataset (Y,X), debiases them and then it needs an estimate of $\sigma$ to compute the p-values for two-sided testing of $\mathcal{H}_{0,j}:\beta_{j}=0$ for $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. Over there, $\widehat{\sigma}$ of $\mathsf{autotune}$ and other benchmarks are supplied and their inference results are compared. We set significance level to 0.05 and if p-value of $j^{th}$ covariate is lesser than 0.05, $\mathcal{H}_{0,j}$ is rejected i.e., it is flagged as a significant predictor and vice versa. Let $\mathcal{S}=\mathopen{}\mathclose{{\left\{j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}:\beta_{j}\neq 0}}\right\}$ be the true support set, then $s=|\mathcal{S}|$ and the false positive rate (FPR) and power becomes
$\mathrm{FPR}$ $=\sum_{j\in\mathcal{S}^{C}}\mathbb{P}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/\mathopen{}\mathclose{{\left(p-s}}\right),~~\mathrm{Power}=\sum_{j\in\mathcal{S}}{\mathbb{P}}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/s.$ (37)
We plot the empirical estimates of these metrics across 100 replications in Fig.˜12, and add a horizontal line $y=0.05$ in the FPR diagram for connecting the empirical FPRs with the significance level used.
@@
Along with Scaled Lasso, we add the Organic and Natural Lasso as [yu2019estimating] focuses on the estimation of error variance in high dimensional regression and are publicly available in the R package $\mathsf{natural}$. Moreover, [yu2019estimating] found organic Lasso to be uniformly better than the CV Lasso based $\widehat{\sigma}$ [reid2016study], hence we skip CV Lasso.
For better understanding of the difficulty in the experiments, we add an oracle method. It has the knowledge of true support set(predictors whose coefficients are nonzero), hence it performs least squares regression of $Y$ on the true support set and report its MSE as the $\widehat{\sigma}^{2}$.
Lastly, in the inference experiment, we drop the results of Organic and Natural Lasso as their $\sigma$ estimation wasn’t competitive in previous experiment.
@@
We focus on the moderate and high-dimensional setups of Tab.˜1, beta types 1˜(B1) and 2˜(B2) and fix the SNR level at 2. We use $\rho=0.35$ to simulate correlated predictors. For tuning in the benchmark methods, the corresponding default settings of their R packages were used.
@@ Results:
We observe in Fig.˜11 that $\mathsf{autotune}$ estimates the $\sigma$ better than all the non-oracle benchmarks. Infact, paired t-test comparisons between $\widehat{\sigma}$’s of $\mathsf{autotune}$ and the benchmarks across 100 replications reveal that the gains in $\sigma$ estimation from $\mathsf{autotune}$ are significant at 0.05 level. The only exception is the setup where the significant predictors are correlated with one another; Beta-type 2 with $\rho=0.35$ where Scaled Lasso matches the performance of $\mathsf{autotune}$. 95% CI of paired t-test comparison between $\mathsf{autotune}$ and Scaled Lasso had 0 in it. Comparison of Fig.˜11(c) and Fig.˜11(d) empirically suggests correlation within the significant predictors hampers the $\widehat{\sigma}$ of $\mathsf{autotune}$ and improves that of Scaled Lasso, but same phenomenon doesn’t happen when significant predictors have correlation primarily with insignificant ones (compare Fig.˜11(a) and Fig.˜11(b)). This observation is in accordance with the comparison of $\mathsf{autotune}$ and Scaled Lasso in terms of RMSE and RTE in Fig.˜5.
The false positive rate(FPR) of all the algorithms was below the significance level of 0.05 across all 100 replications except one. We display side-by-side violin plots of the power of different algorithms across 100 replications, as violin plots give a better comparison of algorithms in terms of power as compared to a jittered scatter plot. The violin plot of $\mathsf{autotune}$’s $\widehat{\sigma}$ closely matches that of oracle’s $\widehat{\sigma}$ across all the setups considered. On the other hand, Scaled Lasso’s $\widehat{\sigma}$ shows suboptimal power everywhere, even in the setup where significant predictors are correlated.
These two experiments support our claim that the partial residuals-driven $\sigma$ estimation mechanism of our $\mathsf{autotune}$ is superior and also exhibits the maximum power empirically possible via $\widehat{\sigma}$ in Lasso inference while controlling the type-1 error rate below the specified level.
@@
Figure 12: Inference of regression coefficients using the estimated $\widehat{\sigma}$ of different algorithms in high dimensional setup
@@ 4.3.2 Visual diagnostics for checking sparsity assumption
We potray the cumulative and adjusted $R^{2}$ of the least squares model formed by incrementally adding predictors to the linear model in the rank order given by $\mathsf{autotune}$ Lasso’s partial residuals. This means in Fig.˜13, a $(s,t)$ point denotes the linear model formed by first $s$ ranked predictors has a cumulative/adjusted $R^{2}$ value of $t$.
@@
We demonstrate the diagnostics for a high dimensional setup of $n=400,p=2000$, SNR =2=2, and predictors are uncorrelated with each other, i.e., $\rho=0$. $s$ is varied between 5 and 10 for the sparse case, 350 for the non-sparse case, and 30 for the semi-sparse case(where it becomes unclear from diagnostics whether sparsity holds or not). We report additional information like the number of iterations of CD $\mathsf{autotune}$ runs to arrive at final $\widehat{\beta}_{\mathsf{autotune}}$, $\mathopen{}\mathclose{{\left\|{\widehat{\beta}_{\mathsf{autotune}}}}}\right\|_{0}$, and estimated support set size (or number of outliers detected by F-tests in the final $\sigma$ update step) at the top of the $R^{2}$ plots.
@@
In Fig.˜13, blue dots refer to predictors with nonzero $\widehat{\beta}$ and red dots means predictors with zero $\widehat{\beta}$. We plot the $R^{2}$ curves till the cumulative $R^{2}$ reaches 1.
@@
The philosophy of our diagnostics is the same as determining the number of principal components from a scree plot. Scientists look for an “elbow” in the scree plot and hope that the principal components appearing before the elbow will give an accurate, sparse (low-dimensional) representation of their high-dimensional data. Similarly, we also try to locate an elbow, a point which is preceded by blue dots with visibly big contributions to the cumulative/adjusted $R^{2}$ and followed by a tight cluster of blue/red dots with marginal contribution to the $R^{2}$. Additionally, the slope of the points will drastically change before and after the elbow spot.
@@
The biggest indicator of sparsity is that the blue dots before the elbow should depict a steep rise in $R^{2}$. Look at Fig.˜13(a), and Fig.˜13(b), where sparsity holds with $s\leq 20$. Sometimes, it’s not clear from the diagnostics whether sparsity holds or not. Consider Fig.˜13(c), where the elbow spot seems to be at the $25^{th}$ predictor. It depicts both the steep rise and slow increase before and after the elbow spot, respectively, but there is not a big change in the slope of the points around the elbow. In Fig.˜13(c), $s$ is 30 and RMSE of $\widehat{\beta}_{\mathsf{autotune}}$ is 0.55. In such experiments where $\mathsf{autotune}$ doesn’t fit well (high RMSE), our diagnostics may not give strong evidence of sparsity.
@@
When sparsity holds, one can gauge the number of important predictors by counting blue dots with visible jumps in the $R^{2}$ plots, just like using scree plot in PCA.
@@
If we are unable to find such an elbow(cumulative R2 plot shows smooth increase) or the elbow spot corresponds to a low cumulative/adjusted R2 (say below 0.5), then we recommend that sparsity doesn’t hold in that dataset. We observe all these negative signs in Fig.˜13(d), which is indeed a non-sparse setup with $s=350$.
Our diagnostics are up to the subjectivity of the user for validating sparsity in Lasso, just like using residual plots to assess whether least square regression gave a good fit to the data. We also highlight that this is the first attempt in the growing Lasso literature to propose an approach to empirically validate the sparsity in high-dimensional data.
@@
5 Financial Application
@@
We take the first $n$ observations in our training dataset, and assign the rest of the observations to the test dataset. We vary $n\in\mathopen{}\mathclose{{\left\{150,160,170,180,190,200,210}}\right\}$ and then compare $\mathsf{autotune}$ against CV Lasso, Scaled Lasso, and Lasso with AIC/BIC tuning in terms of Relative Test Error (RTE), Multiple $\text{R}^{2}$ on test data and sparsity of the final model selected, following the comparison metrics used in [wu2020survey]. Unlike Sec.˜4, here we neither know the true regression coefficients nor the true DGP, so we redefine RTE $\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\widehat{\beta}}}}\right\|_{2}^{2}\big/\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}}}}\right\|_{2}^{2}$. Note that with this definition, we have OOS $^{2}=1-R$TE.
We plot the RTE of the solution path taken by the Lasso using the R function $\mathsf{cv.glmnet}$ (orange) and $\mathsf{autotune}$ (blue) on the test data along with normalized CV errors on the training dataset (red) and demarcate the finally chosen $\lambda$’s with labeled vertical lines in Fig.˜14(a), and 34. We choose this splitting mechanism over others as it ensures test error curve has a “U-shape”.
To further validate $\mathsf{autotune}$’s gains in performance, we generate $B=100$ bootstrap samples of our training set, fit $\mathsf{autotune}$ and benchmarks on each sample, and then evaluate the trained model’s Multiple $\text{R}^{2}$ on the test data, i.e. Out-of-Sample $\text{R}^{2}$ (OOS $\text{R}^{2}$). We report these using side-by-side boxplots in Fig.˜14(b), and 35. AIC and BIC Lasso’s results were omitted as they were not competitive.
In Fig.˜14(a), $\mathsf{autotune}$ quickly reaches it final $\lambda$ within 3 iterations for $n=190$, and $\mathsf{autotune}$’s fitted model also has the lowest RTE. The red training curve shows monotone increase in CV error with increase in $\lambda$ for all $n$, suggesting that CV will lead to overfitting (Fig.˜14(a), 34). Tab.˜3 shows $\mathsf{autotune}$ dominates all the benchmarks across all $n$ in terms of RTE.
For bootstrap results, $\mathsf{autotune}$ shows higher median, 25th, and 75th quantile OOS $\text{R}^{2}$ than all its benchmarks for $n=190$ (Fig.˜14(b)). Bootstrapped OOS R2 gives the same results for other choices of $n$ (see Fig.˜35).
@@
(b) Boxplot of OOS R2 on 100 bootstraped datasets.
Figure 14: 14(a) shows solution paths of Lasso (evaluated by R function $\mathsf{cv.glmnet}$) in orange and $\mathsf{autotune}$ in blue (with endpoint denoted by blue star) for train/test split with $n=190,p=492$. Finally, selected $\lambda$’s are indicated by labeled vertical lines. Fig.˜14(b) gives the boxplot of OOS R2 (=1=1-RTE) of the algorithms trained on 100 bootstraped datasets of the training dataset with $n=190$. The OOS R2 on the original training dataset is indicated by a red diamond point.
We see that $\mathsf{autotune}$ usually gives the sparsest model among all the benchmarks. Scaled Lasso gives the largest estimated model (also observed in [wu2020survey]). All the remaining benchmark methods (excluding CV(1se)) choose the smallest $\lambda$ in the solution path of Lasso (Fig.˜14(a), 34), resulting in overfitted large models.
$n$ $\mathsf{Autotune}$ Scaled CV.min CV.1se AIC BIC

### v2

@@ 4.3 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$
While our main objective in this paper is to design a fast and accurate tuning strategy for Lasso, some intermediate products of our algorithm show promising empirical results in statistical tasks relevant for high-dimensional regression. Here we illustrate some preliminary results for two such tasks, estimation of noise variance $\sigma^{2}$, and a visual diagnostic like scree plot to check the sparsity assumption.
@@ 4.3.1 Noise scale estimation for high-dimensional inference
We empirically contrast $\sigma$ estimation of $\mathsf{autotune}$ against Scaled Lasso, Organic Lasso and Natural Lasso [yu2019estimating]. We also investigate the quality of estimated $\sigma$ of different algorithms by using it in inference of high-dimensional regression coefficients via $\mathsf{hdi}$ R package [dezeure2015high]. We report the sensitivity of $\sigma$ estimation to chosen $\alpha$ in the sequential F-test (see Sigma-Update) of $\mathsf{autotune}$ in the Sec.˜A.5.
@@
We at first compute the difference between estimator of $\sigma^{2}$ of algorithm and the best empirical estimator of noise level, $\sigma^{2}_{\mathrm{empirical}}=\mathrm{Var}(\text{Y}-\text{X}\beta)$ i.e. variance of the residuals obtained via regression of Y on X using true coefficients. Then we plot the boxplots of the difference $\widehat{\sigma}^{2}-\sigma^{2}_{\mathrm{empirical}}$ across 100 replications in Fig.˜10.
For inference, we use De-sparsified Lasso [dezeure2015high] from the $\mathsf{hdi}$ R package. It computes the Lasso coefficients, debiases them, and then it needs an estimate of $\sigma$ to compute the p-values for two-sided testing of $\mathcal{H}_{0,j}:\beta_{j}=0$ for $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. In that stage, $\widehat{\sigma}$ of $\mathsf{autotune}$ and other benchmarks are supplied and their inference results are compared. We set significance level to 0.05 and if p-value of $j^{th}$ covariate is less than 0.05, $\mathcal{H}_{0,j}$ is rejected i.e. it is flagged as a significant predictor and vice versa. Let $\mathcal{S}=\mathopen{}\mathclose{{\left\{j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}:\beta_{j}\neq 0}}\right\}$ be the true support set, then $s=|\mathcal{S}|$ and the false positive rate (FPR) and power becomes
$\mathrm{FPR}$ $=\sum_{j\in\mathcal{S}^{C}}\mathbb{I}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/\mathopen{}\mathclose{{\left(p-s}}\right),~~\mathrm{Power}=\sum_{j\in\mathcal{S}}\mathbb{I}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/s.$ (36)
We plot the empirical estimates of these metrics across 100 replications in Fig.˜11, and add a horizontal line $y=0.05$ in the FPR diagram for connecting the empirical FPR with the significance level used.
@@
Along with Scaled Lasso, we add the Organic and Natural Lasso as [yu2019estimating] focuses on the estimation of error variance in high-dimensional regression and are publicly available in the R package $\mathsf{natural}$. Moreover, [yu2019estimating] found Organic Lasso to be uniformly better than the CV Lasso based $\widehat{\sigma}^{2}$ [reid2016study], hence we skip CV Lasso.
For better understanding of the difficulty in the experiments, we add an oracle method. It has the knowledge of true support set (predictors whose coefficients are nonzero), hence it performs least squares regression of $Y$ on the true support set and report its MSE as the $\widehat{\sigma}^{2}$. Note that, $\sigma^{2}_{\mathrm{empirical}}$ has the knowledge of true $\beta$ whereas $\widehat{\sigma}^{2}$ of the oracle method only has access to the support set of true $\beta$.
Lastly, in the inference experiment, we drop the results of Organic and Natural Lasso as their $\sigma$ estimation was not competitive in the previous experiment of estimating $\sigma$.
@@
We focus on the moderate and high-dimensional setups of Tab.˜1, Beta-types 1˜(B1) and 2˜(B2) and fix the SNR at 2. We use $\rho=0.35$ to simulate correlated predictors. For tuning in the benchmark methods, the corresponding default settings of their R packages were used.
@@ Results:
We observe in Fig.˜10 that $\mathsf{autotune}$ estimates the $\sigma$ better than all the non-oracle benchmarks. In fact, paired t-test comparisons between $\widehat{\sigma}^{2}$ of $\mathsf{autotune}$ and the benchmarks across 100 replications reveal that the gains in $\sigma$ estimation from $\mathsf{autotune}$ are significant at 0.05 level. The only exception is the setup where the significant predictors are correlated with one another; Beta-type 2 with $\rho=0.35$ where Scaled Lasso matches the performance of $\mathsf{autotune}$. 95% CI of paired t-test comparison between $\mathsf{autotune}$ and Scaled Lasso had 0 in it. Comparison of Fig.˜10(c) and Fig.˜10(d) empirically suggests correlation within the significant predictors hampers the $\widehat{\sigma}$ of $\mathsf{autotune}$ and improves that of Scaled Lasso, but same phenomenon does not happen when significant predictors have correlation primarily with insignificant ones (compare Fig.˜10(a) and Fig.˜10(b)). This observation is in accordance with the comparison of $\mathsf{autotune}$ and Scaled Lasso in terms of RMSE and RTE in Fig.˜4.
The false positive rate (FPR) of all the algorithms was below the significance level of 0.05 across all 100 replications except one. We display side-by-side violin plots of the power of different algorithms across 100 replications, as violin plots give a better comparison of algorithms in terms of power as compared to a jittered scatter plot. The violin plot of $\widehat{\sigma}^{2}$ of $\mathsf{autotune}$ closely matches to that of oracle across all the setups considered. On the other hand, $\widehat{\sigma}$ of Scaled Lasso numerically shows suboptimal power everywhere, even in the setup where significant predictors are correlated.
These two experiments support our claim that the PR driven $\sigma$ estimation mechanism of our $\mathsf{autotune}$ is superior to the $\sigma$ estimation mechanism of Scaled Lasso.
@@
Figure 11: Inference of regression coefficients using the estimated $\widehat{\sigma}$ of different algorithms in high-dimensional setup
@@ 4.3.2 Visual diagnostics for checking sparsity assumption
We plot the cumulative and adjusted R2 of the least squares model formed by incrementally adding predictors to the linear model in the rank order given by PR of $\mathsf{autotune}$ Lasso. This means in Fig.˜12, a $(s,t)$ point denotes the linear model formed by first $s$ ranked predictors has a cumulative/adjusted R2 value of $t$.
@@
We demonstrate the diagnostics for a high-dimensional setup of $n=400,p=2000$, SNR =2=2, and predictors are uncorrelated with each other, i.e. $\rho=0$. $s$ is varied between 5 and 10 for the sparse case, 350 for the non-sparse case, and 30 for the semi-sparse case (where it becomes unclear from diagnostics whether sparsity holds or not). We report additional information like the number of iterations of CD $\mathsf{autotune}$ runs to arrive at final $\widehat{\beta}$, $\mathopen{}\mathclose{{\left\|{\widehat{\beta}}}}\right\|_{0}$, and estimated support set size (or number of outliers detected by F-tests in the final $\sigma$ update step) at the top of the R2 plots.
@@
In Fig.˜12, blue dots refer to predictors with nonzero $\widehat{\beta}$ and red dots means predictors with zero $\widehat{\beta}$. We plot the R2 curves till the cumulative R2 reaches 1.
@@
The intuition behind our diagnostics is the same as determining the number of principal components from a scree plot. Scientists look for an “elbow” in the scree plot and hope that the principal components appearing before the elbow will give an accurate, sparse (low-dimensional) representation of their high-dimensional data. Similarly, we also try to locate an elbow, a point which is preceded by blue dots with visibly big contributions to the cumulative or adjusted R2 and followed by a tight cluster of blue and red dots with marginal contribution to the R2. Additionally, the slope of the points will drastically change before and after the elbow spot.
@@
The biggest indicator of sparsity is that the blue dots before the elbow should depict a steep rise in R2. Note that in Fig.˜12(a), and Fig.˜12(b), where sparsity holds with $s\leq 20$. Sometimes, it is not clear from the diagnostics whether sparsity holds or not. Consider Fig.˜12(c), where the elbow spot seems to be at the $25^{th}$ predictor. It depicts both the steep rise and slow increase before and after the elbow spot, respectively, but there is not a big change in the slope of the points around the elbow. In Fig.˜12(c), $s$ is 30 and RMSE of $\widehat{\beta}_{\mathsf{autotune}}$ is 0.55. In such experiments where $\mathsf{autotune}$ does not fit well (high RMSE), our diagnostics may not give strong evidence of sparsity.
@@
When sparsity holds, one can gauge the number of important predictors by counting blue dots with visible jumps in the R2 plots, just like using scree plot in PCA.
@@
If we are unable to find such an elbow (cumulative R2 plot shows smooth increase) or the elbow spot corresponds to a low cumulative or adjusted R2 (say below 0.5), then we recommend that sparsity doesn’t hold in that dataset. We observe all these negative signs in Fig.˜12(d), which is indeed a non-sparse setup with $s=350$.
@@
5 Empirical illustration on a financial dataset
@@
We take the first $n$ observations in our training dataset, and assign the rest of the observations to the test dataset. We vary $n\in\mathopen{}\mathclose{{\left\{150,160,170,180,190,200,210}}\right\}$ and then compare $\mathsf{autotune}$ against CV Lasso, Scaled Lasso, and Lasso with AIC and BIC tuning in terms of Relative Test Error (RTE), multiple $\text{R}^{2}$ on test data and sparsity of the final model selected, following the comparison metrics used in [wu2020survey]. We choose this splitting mechanism over others as it ensures test error curve of CV Lasso has a “U-shape”. Unlike Sec.˜4, here we neither know the true regression coefficients nor the true DGP, so we redefine RTE $\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\widehat{\beta}}}}\right\|_{2}^{2}\big/\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}}}}\right\|_{2}^{2}$. We define multiple R2 on test data as the out-of-sample R2 (O.O.S. R2). Note that with these two definitions, we have O.O.S. $^{2}=1-R$TE.
We plot the RTE of the solution path taken by the Lasso using the R function $\mathsf{cv.glmnet}$ (orange) and $\mathsf{autotune}$ (blue) on the test data along with normalized CV errors on the training dataset (red) and demarcate the finally chosen $\lambda$’s with labeled vertical lines in Fig.˜13(a), and 33. To further validate $\mathsf{autotune}$’s gains in performance, we generate $B=100$ bootstrap samples of our training set, fit $\mathsf{autotune}$ and benchmarks on each sample, and then evaluate the trained model’s Multiple $\text{R}^{2}$ on the test data, i.e. Out-of-Sample $\text{R}^{2}$ (O.O.S. $\text{R}^{2}$). We report these using side-by-side boxplots in Fig.˜13(b), and 34. Results of AIC and BIC Lasso were omitted as they were not competitive.
In Fig.˜13(a), $\mathsf{autotune}$ quickly reaches it final $\lambda$ within 3 iterations for $n=190$, and $\mathsf{autotune}$’s fitted model also has the lowest RTE. The red training curve shows monotone increase in CV error with increase in $\lambda$ for all $n$, suggesting that CV will lead to overfitting (Fig.˜13(a), 33). Tab.˜3 shows $\mathsf{autotune}$ dominates all the benchmarks across all $n$ in terms of RTE.
For bootstrap results, $\mathsf{autotune}$ shows higher median, 25th, and 75th quantile O.O.S. $\text{R}^{2}$ than all its benchmarks for $n=190$ (Fig.˜13(b)). Bootstrapped O.O.S. R2 gives the same results for other choices of $n$ (see Fig.˜34).
@@
(b) Boxplot of O.O.S. R2 on 100 bootstraped datasets.
Figure 13: 13(a) shows solution paths of Lasso (evaluated by R function $\mathsf{cv.glmnet}$) in orange and $\mathsf{autotune}$ in blue (with endpoint denoted by blue star) on test datafold with $n_{\mathrm{train}}=190,n_{\mathrm{test}}=62,p=492$. Finally, selected values of $\lambda$ are indicated by labeled vertical lines. Fig.˜13(b) gives the boxplot of O.O.S. R2 (=1=1-RTE) of the algorithms trained on 100 bootstraped datasets of the training dataset with $n=190$. The O.O.S. R2 on the original training dataset is indicated by a red diamond point.
We see that $\mathsf{autotune}$ usually gives the sparsest model among all the benchmarks. Scaled Lasso gives the largest estimated model (also observed in [wu2020survey]). All the remaining benchmark methods (excluding CV(1se)) choose the smallest $\lambda$ in the solution path of Lasso (Fig.˜13(a), 33), resulting in overfitted large models.
$n$ $\mathsf{Autotune}$ Scaled CV(min) CV(1se) AIC BIC